# The Impact of Surface Co-location and Eye-tracking on Mixed Reality Typing

### v2

@@ Abstract.
Accuracy and speed are pivotal when typing. We hypothesized that the lack of tactile feedback on midair mixed reality keyboards may adversely impact performance. Our first experiment investigated providing tactile feedback to users typing in mixed reality by co-locating the virtual keyboard on a table or a wall. The keyboard was deterministic (without auto-correct), allowed mixed case typing with symbols, and relied only on the hand-tracking provided by a commodity headset’s egocentric cameras. Users preferred and had the highest entry rate of 12 words-per-minute using a midair keyboard. Error rates were similar in all conditions. Based on user feedback, our second experiment explored ten-finger typing. We used a novel eye-tracking technique to mitigate accidental key presses. This technique was preferred for ten-finger typing and halved the number of times backspace was pressed. However, participants were faster using only their index fingers without eye-tracking at 11 words-per-minute.
@@ 1. Introduction
Mixed reality (MR) offers the ability to project virtual objects onto the real world and interact with them. As MR technology develops, it may begin to displace other portable computing devices such as cell phones and laptops. Users might wish to utilize MR to perform everyday tasks. Typing is critical to efficiently performing common tasks such as writing emails, browsing the web, and creating documents. Unlike a traditional keyboard where you can feel the locations of the keys and when they are actuated, MR systems often employ midair keyboards which lack tactile feedback. The loss of tactile feedback can impair users’ accuracy when typing (Rabin and Gordon, 2004), potentially reducing the utility of MR compared to conventional devices for supporting common, everyday computing activities.
@@
Speech input offers an alternative to typing but may not always be socially acceptable (Baier and Burmester, 2019), and recognition of difficult words can be a problem (Adhikary and Vertanen, 2021). Auxiliary devices such as physical keyboards (McGill et al., 2015; Walker et al., 2017; Knierim et al., 2018; Grubert et al., 2018), controllers (Speicher et al., 2018a; Xu et al., 2019a), gloves (Markussen et al., 2014; McCaul and Sutherland, 2004; He et al., 2022), or watches (Ahn et al., 2017), could help but require carrying additional devices. Other solutions re-imagine the text input process (e.g. word-gesture keyboards (Kristensson and Zhai, 2004; Zhai and Kristensson, 2012; Ahn et al., 2017), dwell-free gaze typing (Kristensson and Vertanen, 2012; Hu et al., 2024), or finger gestures (McCaul and Sutherland, 2004; He et al., 2022; Feit et al., 2015)). However, such methods introduce a learning curve that may be a barrier to MR adoption. Compared to past work, we focus on enabling mixed reality typing without: 1) requiring users to learn a new interaction technique, 2) requiring devices or sensors aside from the headset, and 3) using a language model to compensate for noisy input.
@@
We conjectured providing tactile feedback would improve typing in MR. In Experiment 1, we aligned a virtual keyboard with a wall or table. This adds tactile feedback to keypresses and could be used whenever a flat surface is present. Although surface co-location may not be feasible in all situations, people are frequently in human-made environments where flat surfaces are prevalent. We compared the surface aligned keyboards with one floating in midair. Participants typed using only their index fingers. We found participants preferred the midair keyboard compared to keyboards near a table or wall. They typed faster in midair at 12.2 words-per-minute (WPM) compared to on a table (9.1 WPM) or a wall (8.8 WPM). Experiment 2 focused on ten-finger typing, a feature requested by many participants in our first study. Pilot testing revealed ten-finger typing in midair led to numerous accidental keystrokes. Therefore, in addition to comparing ten- and index-finger typing, we also added a novel eye-tracking feature designed to prevent accidental keystrokes. Our feature filtered out key presses not located near a user’s eye gaze location. We theorized since users were likely visually guiding their typing due to the lack of tactile feedback, this feature would not be an undue burden. We found users preferred index- to ten-finger typing. We found eye-tracking reduced accidental key presses, improved accuracy, and slightly increased entry rates for ten-finger typing (7.4 WPM versus 7.1 WPM).
@@
User feedback was in general positive for the tactile feedback and ergonomics offered by virtual keyboards located near a surface. However near surfaces, accurate hand-tracking was a challenge for a current commercial headset.
@@
We release detailed logs from our studies and the source code of our HoloLens keyboard experimental applications.
@@ 2. Related Work
If mixed reality is going to be used to perform everyday computing tasks, text input needs to be efficient. Many solutions have been proposed to address this need. A midair virtual keyboard that allows users to type directly on it with their hands is both familiar and easy-to-learn. Early systems such as ARKB (Lee and Woo, 2003) relied on external cameras to track a user’s hands over a virtual keyboard. Markussen et al. (Markussen et al., 2013) used an array of 24 external cameras to track a user’s hands. Users directly selected keys on a projected QWERTY keyboard, or used ambiguous gestures to specify keys. Direct typing was the fastest at 13 WPM after six sessions. Dudley et al. (Dudley et al., 2018) investigated midair virtual keyboard typing on a HoloLens 1. Similar to our work, they relied only on the built-in headset sensors. Despite the HoloLens 1 only tracking a user’s wrist position, users typed at 18 WPM on an auto-correcting keyboard with word predictions. In comparison, we used a HoloLens 2 with full hand-tracking and focused on input of more difficult mixed case text with symbols and numbers.
Cheng et al. (Cheng et al., 2022) explored if sitting at a table improved users’ ability to select and drag targets in virtual reality (VR). Users’ arms and finger were tracked via external cameras. They found the presence of a table improved user accuracy and reduced fatigue of horizontal and vertical midair interfaces. In this work, we also chose to use a table as it is a common everyday object. However, we focus on MR and the more accuracy demanding task of keyboard typing.
@@
Dudley et al. (Dudley et al., 2019) had users type with two or ten fingers on virtual keyboards located near a surface or in midair. Users’ hands were tracked via an array of external cameras and markers on the hands. The system used an auto-correct algorithm based on the system knowing the target text. Users typed significantly faster on the surface keyboards with either two fingers (56 WPM versus 42 WPM) or all ten fingers (52 WPM versus 35 WPM). Richardson et al. (Richardson et al., 2020) also investigated surface typing by tracking users’ hands via external cameras. Users typed phrases quickly at 73 WPM. Using a neural network hand motion model and a language model, they found recorded typing sequences could be decoded at a low error rate of 2%. In recent followup work, Richardson et al. (Richardson et al., 2024) used only the egocentric cameras of a VR headset to infer keypresses on a virtual keyboard situated on a desk. Users typed at 44 WPM with an uncorrected error rate of 7%. While the system had no explicit language model, the motion model may have learned one as the majority of its training data was English text. Indeed, on non-language text, offline experiments showed a much high error rate of 16% compared to 5% on English. Similar to (Richardson et al., 2024), we utilize the hand-tracking available from the egocentric cameras of a commodity headset. However, we rely solely on a human-in-the-loop process where the user observes their inferred finger locations and uses this feedback to supervise their typing. A system free from language-based prediction bias may be needed in some scenarios (e.g. typing a password).
@@
Moving somewhat away from the metaphor of tapping on a keyboard, a word-gesture keyboard (Kristensson and Zhai, 2004; Zhai and Kristensson, 2012) allows writing by swiping through a word’s letters. The Vulture keyboard (Markussen et al., 2014) allowed writing on a keyboard using a glove and an external camera array. After ten sessions, users wrote at 21 WPM. Moving further away from conventional typing, an algorithm can infer a user’s intended text from ambiguous gestures. Early work allowed users wearing a glove to write words in VR by selecting different sets of letters by flexing their finger (McCaul and Sutherland, 2004). A similar recent system is TapGazer (He et al., 2022) which uses a glove to detect which finger a user is tapping with disambiguation via eye gaze or additional taps.
@@
Other approaches for text input in VR and AR have explored input methods other than hand gestures. The direction of a user’s head can be used to select a letter from a circular menu (Xu et al., 2019b) or a QWERTY keyboard (Yu et al., 2017; Speicher et al., 2018a; Gizatdinova et al., 2018; Xu et al., 2019). While Gizatdinova et al. (Gizatdinova et al., 2018) found head-based typing was effective on small keyboards, users thought it felt unnatural and they preferred gaze-based typing. Gaze-based typing in VR has been investigated both as the sole input method (Lu et al., 2021) and in combination with head gestures (Feng et al., 2021). However, the use of dwell time, the minimum time a user must look at a key in order to trigger it, can lead to eyestrain (Lystbæk et al., 2022; Pfeuffer et al., 2020). An alternate approach, dwell-free eye typing (Kristensson and Vertanen, 2012), allows users to gaze continually through letters. Using this approach, Hu et al. (Hu et al., 2024) found users wrote at 12 WPM using a HoloLens 2 headset. While potentially fast, dwell-free typing requires user learning and may not work for hard-to-predict text.
@@
Using speech input can help reduce MR challenges such as robustly sensing users hands and accurately displaying visual feedback. In early work, Bowman et al. (Bowman et al., 2002) explored VR text input using Wizard of Oz speech recognition and input via a glove, keyboard, or tablet. Speech recognition was fastest at around 13 WPM. In SWIFTER (Pick et al., 2016), users wrote in a virtual environment at 24 WPM using speech with correction via a hand controller. Similar to our system, Adhikary and Vertanen (Adhikary and Vertanen, 2021) eschewed the use of auxiliary devices and relied solely on speech and hand-tracking for VR text input. Speaking and then correcting via hand gestures was the fastest at 28 WPM. Notably, without speech, users wrote at 11 WPM on a midair auto-correcting keyboard with word predictions. This entry rate was similar to our users’ midair deterministic keyboard performance. While speech input can be a good option, it may not always be socially acceptable (Baier and Burmester, 2019), and accuracy can be poor in noisy environments or for challenging text (e.g. usernames and passwords).
@@
Most existing MR typing solutions, while often allowing typing with both hands, typically only support typing with the index fingers (Lystbæk et al., 2022; Dudley et al., 2018; Adhikary and Vertanen, 2021). Systems that do allow use of all fingers generally require an auxiliary input device (He et al., 2022), external sensing device (Yi et al., 2015), or physical keyboard (McGill et al., 2015; Walker et al., 2017; Knierim et al., 2018; Grubert et al., 2018). Ten-finger typing is challenging as the dexterity limitations of the human hand can make some multi-finger gestures difficult to perform without the tactile assistance of a physical keyboard (Sridhar et al., 2015). Additionally, ten-finger typing on virtual keyboards can lead to accidental inputs and key coactivations (Foy et al., 2021). While multi-finger gestures have been used for navigating MR interfaces, the small movements required for typing may be difficult for vision-based hand-tracking systems to recognize (Malik et al., 2005). To try and improve ten-finger typing, in Experiment 2, we use eye-tracking to filter out key presses not located near a user’s gaze. Past work shows that users will naturally tend to look at the keys they are pressing (Lystbæk et al., 2022).
@@ 3. Experiment 1
In our first experiment, participants wore a HoloLens 2 headset and typed sentences on a virtual keyboard using their index fingers. We measured user performance and preference for typing on a midair keyboard and keyboards co-located with a table and with a wall. Our primary goal was to assess the impact of the tactile feedback offered by the surface co-location. We opted not to include a physical keyboard condition since typing on physical keyboards has been well-studied and our focus is on on-the-go applications where all sensing is performed by the headset. Since midair typing is currently a common choice for MR typing applications, we used a midair keyboard as a baseline.
@@ 3.1. Design
Table — The keyboard was placed just above a table, using a QR code on the table to ensure consistent spatial placement. The keyboard was placed horizontally, such that the back of the keyboard was parallel with the table.
Wall — The keyboard was placed slightly in front of a wall, using the same QR code as a spatial anchor. The keyboard was placed vertically, such that the back of the keyboard was parallel with the wall.
We hypothesized the tactile feedback conditions (i.e. Wall and Table) would perform better than the non-tactile feedback condition (i.e. Midair). We also hypothesized Table would be preferred due to its similarity to a standard physical keyboard. We investigated both the Wall and Table locations to assess typing on common real-world flat surfaces. In addition, the keyboard position and orientation relative to the headset is different between the Wall and Table conditions. This allowed us to investigate whether this impacted hand-tracking accuracy.
In all conditions, participants used their index fingers on one or both hands to enter text. This choice prioritizes ecological validity by allowing participants to type as they would in the real world. Participants also could choose to rest their arms on a table as needed to potentially reduce fatigue. Our keyboard was deterministic and did not auto-correct or predict words. We did this to understand the raw performance of typing as might be required for text not easily predicted by a language model.
All conditions had audio and visual feedback for key presses. A key was triggered when either index finger entered the key’s bounding area, pushed the key’s face a certain depth, and exited the same key’s bounding area. On a key press, we played a button sound effect and displayed a white border around the key. We displayed participants’ hands using the Unity MR toolkit’s default hand skin, an overlay of white and grey triangles. We displayed blue spheres on fingertips and grey cubes on hand joints (Figure 1 right). The keyboard was 44 cm 16 cm and rendered slightly transparent blue with white key labels. We used a full QWERTY keyboard including numbers, symbols, shift keys, caps lock, and backspace. Users could see their current text above the keyboard and could correct errors via a backspace key.

### v3

# Exploring Keyboard Positioning and Ten-Finger Typing in Mixed Reality

@@ Abstract.
Accuracy and speed are pivotal when typing. Mixed reality typing is typically performed by typing on a midair keyboard with your index fingers. This deprives users of both the tactile feedback available on physical devices and the ability to press keys with the most convenient finger. Our first experiment investigated providing tactile feedback by positioning the virtual keyboard on a table or wall. The keyboard was deterministic (without auto-correct), supported mixed case typing with symbols, and relied only on the hand-tracking provided by a commodity headset’s egocentric cameras. Users preferred and had the highest entry rate of 12 words-per-minute using a midair keyboard. Error rates were similar in all conditions. Our second experiment explored ten-finger typing and used a novel eye-tracking technique to avoid accidental key presses. This technique was preferred for ten-finger typing and halved corrections. However, participants were faster using their index fingers without eye-tracking at 11 words-per-minute.
@@ 1. Introduction
Mixed reality (MR) offers the ability to project virtual objects onto the real world and interact with them. As MR technology develops, users might wish to utilize MR to perform everyday tasks. Typing is critical to efficiently performing common tasks such as writing emails, browsing the web, and creating documents. Unlike a traditional keyboard where you can feel the locations of the keys and when they are actuated, MR systems often employ midair keyboards which lack tactile feedback. The loss of tactile feedback has been shown to impair users’ accuracy when typing (Rabin and Gordon, 2004). Further, MR systems typically restrict typing to just a user’s index fingers. These issues can potentially reduce the utility of MR compared to conventional devices for supporting common, everyday computing activities.
@@
Speech input offers an alternative to typing but may not always be socially acceptable (Baier and Burmester, 2019), and recognition of difficult words can be a problem (Adhikary and Vertanen, 2021a). Auxiliary devices such as physical keyboards (McGill et al., 2015; Walker et al., 2017; Knierim et al., 2018; Grubert et al., 2018), controllers (Speicher et al., 2018a; Xu et al., 2019a), gloves (Markussen et al., 2014; McCaul and Sutherland, 2004; He et al., 2022), or watches (Ahn et al., 2017), could help but require carrying additional devices. Other solutions re-imagine the text input process (e.g. word-gesture keyboards (Kristensson and Zhai, 2004; Zhai and Kristensson, 2012; Ahn et al., 2017), dwell-free gaze typing (Kristensson and Vertanen, 2012; Hu et al., 2024), or finger gestures (McCaul and Sutherland, 2004; He et al., 2022; Feit et al., 2015)). However, such methods introduce a learning curve that may be a barrier to MR adoption. Such methods also usually rely on a language model which restricts input to predictable natural language text. Compared to past work, we focus on mixed reality typing that: 1) leverages users’ existing knowledge of tapping on an virtual keyboard, 2) uses only the sensors built into the headset, and 3) allows input of arbitrary character sequences (i.e. without relying on a language model to compensate for noisy input).
Tactile feedback of a virtual keyboard on a table has been shown to improve performance (Dudley et al., 2019). What hasn’t been shown is whether this is feasible using only the tracking available from the egocentric sensors on the headset and without relying on auto-correct. Further, only a keyboard positioned on a horizontal surface has been explored previously.
@@
In Experiment 1, we explored three keyboard positions: midair, table aligned, and wall aligned. The table and wall surfaces add tactile feedback to keypresses and could be used whenever a flat surface is present. Although locating a keyboard on a surface may not be feasible in all situations, people are frequently in human-made environments where flat surfaces are prevalent. Participants typed using only their index fingers. They typed faster in midair at 12.2 words-per-minute (WPM) compared to on a table (9.1 WPM) or a wall (8.8 WPM). We found participants preferred the midair keyboard compared to keyboards near a table or wall.
Experiment 2 focused on ten-finger typing, a feature requested by many participants in our first study. Pilot testing revealed ten-finger typing in midair led to numerous accidental keystrokes. Therefore, in addition to comparing ten- and index-finger typing, we also added a novel eye-tracking feature designed to prevent accidental keystrokes. Our feature filtered out key presses not located near a user’s eye gaze location. We theorized since users were likely visually guiding their typing due to the lack of tactile feedback, this feature would not be an undue burden. We found users preferred index- to ten-finger typing. We found eye-tracking reduced accidental key presses, improved accuracy, and slightly increased entry rates for ten-finger typing (7.4 WPM versus 7.1 WPM).
@@
While many users liked the tactile feedback afforded by a keyboard on a table or wall, we show midair typing provides higher performance on a current commercial headset due to challenges of hand-tracking near a surface.
@@
We release detailed logs from our studies and the source code of our HoloLens111HoloLens is an augmented reality headset released by Microsoft in 2016. A second version of the headset (and the one we used) was released in 2019. experimental applications.
@@ 2. Related Work
2.1. Midair virtual keyboards
For mixed reality to be useful for everyday computing tasks, text input needs to be efficient. Many solutions have been proposed to address this need. A midair virtual keyboard allowing users to type directly with their hands is both familiar and easy-to-learn. Early systems such as ARKB (Lee and Woo, 2003) relied on external cameras to track a user’s hands over the keyboard. Similar to our first study, ARKB used printed fiducial markers to position a keyboard horizontally and slightly above a table. No user study results were reported. Markussen et al. (2013) also relied on external cameras for hand-tracking. Users either directly selected keys or used ambiguous gestures to specify keys. The keyboard was projected vertically in front of users on a large display. Direct typing was the fastest at 13 WPM after six sessions.
Dudley et al. (2018) investigated midair typing on a HoloLens 1. The keyboard was located in midair slightly below a user’s eyes and was tilted 20 degrees away from vertical. Similar to our work, they relied only on the built-in headset sensors. Despite the HoloLens 1 only tracking a user’s wrist position, users typed at 18 WPM on an auto-correcting keyboard with word predictions. In comparison to past work, we used a HoloLens 2 with full hand-tracking, eschewed the use of external cameras, and focused on the input of more difficult mixed case text with symbols and numbers.
2.2. Physical keyboards in VR and AR
Various work has explored typing on a physical keyboard while wearing a VR or AR headset. McGill et al. (2015) had users type: 1) without a headset, 2) wearing a headset with no view of the keyboard or a user’s hands, 3) wearing a headset with a view of the keyboard and a user’s hands blended in, or 4) wearing a headset with a view of the real-world. Users were fastest without the headset (59 WPM) and slowest when they had no view of the keyboard or their hands (24 WPM). Blending or full real-world visual feedback provided similar performance at 39 and 37 WPM respectively.
Rather than blending real-world video into VR, Walker et al. (2017) highlighted keys on a virtual keyboard as a user typed. Coupled with auto-correct, users typed in VR with only a virtual keyboard at 44 WPM. Chiossi et al. (2024) had users type in AR, VR, and augmented virtuality (where real-world imagery is displayed in VR (McGill et al., 2015)). Users typed at similar speeds of 34.9, 32.5, and 34.5 WPM respectively. While typing on a physical keyboard can be fast and support rich text input, users may not want to carry an extra device solely for text entry. As such, we focus on using a virtual keyboard.
2.3. Tactile feedback
@@
Dudley et al. (2019) had users type with two or ten fingers on virtual keyboards located near a surface or in midair. Users’ hands were tracked via an external camera array and markers on their hands. Users were seated at a desk and the midair keyboard was tilted slightly above horizontal. The system used an auto-correct algorithm based on the known target text. Users typed faster on the surface keyboards with either two fingers (56 WPM versus 42 WPM) or all ten fingers (52 WPM versus 35 WPM). Our work focuses on input without auto-correct and without knowing the target text.
@@
Richardson et al. (2020) also investigated surface typing by tracking users’ hands via external cameras. Users typed phrases quickly at 73 WPM. Using a neural network hand motion model and a language model, they typed sequences could be decoded at a low error rate of 2%. In followup work, Richardson et al. (2024) used only the egocentric cameras of a VR headset to infer keypresses on a virtual keyboard situated on a desk. Users typed at 44 WPM with an uncorrected error rate of 7%. While the system had no explicit language model, the motion model may have learned one as the majority of its training data was English text. Indeed, on non-language text, offline experiments showed a much high error rate of 16% compared to 5% on English. Similar to Richardson et al. (2024), we utilize the hand-tracking available from the egocentric cameras of a commodity headset. However, we rely solely on a human-in-the-loop process where the user observes their inferred finger locations and uses this feedback to supervise their typing. A system free from language-based prediction bias may be needed in some scenarios (e.g. typing a password).
2.4. Interaction context
Mixed reality offers the possibly of interaction in many real-world contexts, for example while sitting, standing, or walking. Cheng et al. (2022) explored if sitting at a table improved users’ ability to select and drag targets in VR. Users’ arms and finger were tracked via external cameras. They found having a table improved user accuracy and reduced fatigue of horizontal and vertical midair interfaces. Dudley et al. (2019) controlled for posture by having users sit at a desk while they typed on a virtual keyboard. We also chose to investigate seated interaction as this is a common context in which people do productivity related text entry. Further, given our chosen task of entering rich text without auto-correct, we wanted to reduce physical fatigue and user motion that could make hand-tracking interaction more challenging.
2.5. Alternatives to QWERTY typing
@@
Moving somewhat away from the metaphor of tapping on a keyboard, a word-gesture keyboard (Kristensson and Zhai, 2004; Zhai and Kristensson, 2012) allows writing by swiping through a word’s letters. The Vulture keyboard (Markussen et al., 2014) allowed writing on a keyboard using a glove and an external camera array. After ten sessions, users wrote at 21 WPM. Moving further away from conventional typing, an algorithm can infer a user’s intended text from ambiguous gestures. Early work allowed users wearing a glove to write words in VR by selecting different sets of letters by flexing their finger (McCaul and Sutherland, 2004). A similar recent system is TapGazer (He et al., 2022) which uses a glove to detect finger taps with disambiguation via eye gaze or additional taps.
@@
Other approaches for text input in VR and AR have explored input methods other than hand gestures. The direction of a user’s head can be used to select a letter from a circular menu (Xu et al., 2019b) or a QWERTY keyboard (Yu et al., 2017; Speicher et al., 2018a; Gizatdinova et al., 2018; Xu et al., 2019). While Gizatdinova et al. (2018) found head-based typing was effective on small keyboards, users thought it felt unnatural and they preferred gaze-based typing. Gaze-based typing in VR has been investigated both as the sole input method (Lu et al., 2021) and in combination with head gestures (Feng et al., 2021). However, the use of dwell time, the minimum time a user must look at a key in order to trigger it, can lead to eyestrain (Lystbæk et al., 2022; Pfeuffer et al., 2020). An alternate approach, dwell-free eye typing (Kristensson and Vertanen, 2012), allows users to gaze continually through letters. Using this approach, Hu et al. (2024) found users wrote at 12 WPM using a HoloLens 2 headset. While potentially fast, dwell-free typing requires user learning and may not work for hard-to-predict text.
2.6. Speech input
@@
Using speech input can help reduce MR challenges such as robustly sensing users hands and accurately displaying visual feedback. In early work, Bowman et al. (2002) explored VR text input using Wizard of Oz speech recognition and input via a glove, keyboard, or tablet. Speech recognition was fastest at around 13 WPM. In SWIFTER (Pick et al., 2016), users wrote in a virtual environment at 24 WPM using speech with correction via a hand controller. Similar to our system, Adhikary and Vertanen (2021a) avoided auxiliary devices and relied solely on speech and hand-tracking for VR text input. Speaking and then correcting via hand gestures was the fastest at 28 WPM. Notably, without speech, users wrote at 11 WPM on a midair auto-correcting keyboard with word predictions. This entry rate was similar to our users’ midair deterministic keyboard performance. While speech input can be a good option, it may not always be socially acceptable (Baier and Burmester, 2019), and accuracy can be poor in noisy environments or for challenging text (e.g. usernames and passwords).
2.7. Ten finger interaction
@@
While existing MR systems often allow typing with both hands, typically they only support use of the index fingers (Lystbæk et al., 2022; Dudley et al., 2018; Adhikary and Vertanen, 2021a). Systems that allow use of all fingers generally require an auxiliary input device (He et al., 2022), external sensing device (Yi et al., 2015), or physical keyboard (McGill et al., 2015; Walker et al., 2017; Knierim et al., 2018; Grubert et al., 2018). Ten-finger typing is challenging as the dexterity limitations of the human hand can make some multi-finger gestures difficult without the tactile assistance of a physical keyboard (Sridhar et al., 2015). Additionally, ten-finger typing on virtual keyboards can lead to accidental key coactivations (Foy et al., 2021). While multi-finger gestures have been used for navigating MR interfaces, the small movements required for typing may be difficult for vision-based hand-tracking systems to recognize (Malik et al., 2005). Past work has showed that users naturally look at the keys they are pressing (Lystbæk et al., 2022). Thus in Experiment 2, we explore using eye-tracking to filter out key presses not near a user’s gaze.
@@ 3. Experiment 1
In our first experiment, participants wore a HoloLens 2 headset and typed sentences on a virtual keyboard using their index fingers. We measured user performance and preference for typing on a midair keyboard and keyboards located on a table and on a wall. Our primary goal was to assess the impact of different keyboard positions, including placement on a horizontal and vertical surface. We opted not to include a physical keyboard condition since typing on physical keyboards has been well-studied and our focus is on on-the-go applications where all sensing is performed by the headset. Since midair typing is currently a common choice for MR typing applications, we used a midair keyboard as a baseline.
@@ 3.1. Design
Midair — The keyboard was placed in midair in front of the participant. The keyboard was horizontally centered above the table with the bottom of the spacebar key being 19.0 cm from the front of the table and 4.5 cm above the table. This position made it possible for users to rest their forearms against the edge of the table while typing if they so desired. The keyboard was angled 15 degrees above horizontal. In pilot testing, we found this tilt made it more comfortable to type while seated compared to a vertical or completely horizontal keyboard. A similar tilt was used in a study where participants typed on or above a surface while seated (Dudley et al., 2019) (c.f. Figure 5).
Table — The keyboard was placed just above a table. The keyboard was horizontally centered on the table with the bottom of the spacebar key being 28.5 cm from the front of the table. This position made it possible for users to rest their arms on the table as they typed if they so desired.
Wall — The keyboard was placed slightly in front of a wall. The keyboard was horizontally centered on the table with the bottom of the spacebar being 16.0 cm above the table. This position coupled with the small size of the table made it possible for users to rest their arms on the table as they typed if they so desired.
We hypothesized the keyboards positions on surfaces (i.e. Wall and Table) would perform better than the baseline condition (i.e. Midair). We also hypothesized Table would be preferred due to its similarity to a standard physical keyboard. We investigated both the Wall and Table locations to assess typing on common real-world flat surfaces. In addition, the keyboard position and orientation relative to the headset is different between the Wall and Table conditions. This allowed us to investigate whether this impacted hand-tracking accuracy.
In all conditions, participants typed with their index fingers on one or both hands. This choice prioritizes ecological validity by allowing participants to type as they would in the real world. Our keyboard was deterministic and did not perform auto-correction or predict words. We did this to understand the raw performance of typing as might be required for text not easily predicted by a language model.
All conditions had audio and visual feedback for key presses. A key was triggered when either index finger entered the key’s bounding area, pushed the key’s face a certain depth, and exited the same key’s bounding area. On a key press, we played a sound effect and displayed a white border around the key. We displayed participants’ hands using the Unity MR toolkit’s default hand skin, an overlay of white and grey triangles. We displayed blue spheres on index fingertips and grey cubes at joints.
The keyboard size (including four character rows plus the spacebar) was 33.0 cm 11.0 cm. Each letter key was 2.2 cm 2.2 cm. The keyboard was similar in size to a physical desktop keyboard. It is also similar to the virtual key sizes reported in past work, e.g. 2.2 cm (Dudley et al., 2018), 2.5 cm (Dudley et al., 2019), and 1.9 cm (Richardson et al., 2024). The keyboard was rendered slightly transparent blue with white key labels. We used a full QWERTY keyboard including numbers, symbols, shift keys, caps lock, and backspace. Users could see their current text above the keyboard and could correct errors via a backspace key.